# KnobGen: Controlling the Sophistication of Artwork in Sketch-Based Diffusion Models

### v2

@@ 3.2.1 Macro Pathway
Diffusion models typically rely on text-based conditioning using CLIP text encoders (Radford et al., 2021) to capture high-level semantics (Ramesh et al., 2021; Saharia et al., 2022; Nichol et al., 2021), but this approach often misses out on structural cues inherent to other modalities, such as sketches. Although models such as CLIP (Radford et al., 2021) encode visual features and textual semantics, they remain biased toward coarse-grained features (Bianchi et al., 2024; Wang et al., 2023). In our CGC module, Figure 5.B, we used this fact to our advantage to fuse a high-level visual and linguistic understanding to control DM generation by incorporating both text and image embeddings through a cross-attention mechanisms.
@@
In our CGC module, we leverage the trained CLIP text encoder and its corresponding image encoder variant available in the pretrained Stable Diffusion Model (Rombach et al., 2022). Our CGC module first takes the raw sketch image (condition) and prompt as input. Using the CLIP image and text encoders, the CGC module first projects them into $x_{i}\in\mathbb{R}^{256\times 1024}$ and $x_{p}\in\mathbb{R}^{77\times 768}$ which are the image and text embeddings. A cross-attention mechanism then fuses these embeddings to produce a multimodal representation that combines textual semantics and visual cues, Figure 5.B. This enables the diffusion process to encode the high-level semantics from text while explicitly integrating spatial features from the sketch using the Clip image encoder. The cross-attended embeddings are injected into layers of the denoising U-Net to preserve the coarse-grained visual-textual features throughout the diffusion process. Detailed discussion of the CGC module can be found in the Appendix B.
@@ 3.2.2 Micro Pathway
For artistic users, preserving fine-grained details such as object boundaries and textures is essential. The Fine-Grained Controller (FGC) is designed to address these requirements by integrating pretrained modules such as ControlNet (Zhang et al., 2023) and the T2I-Adapter (Mou et al., 2024), which excel in capturing these intricate features. Our Micro Pathway can utilize any pretrained fine-grained controller module which shows the flexibility of our proposed framework.
@@ 3.3 Modulator at training
One of the key innovations in KnobGen is the tanh-based modulator, which regulates the contributions of the micro and macro pathways during training, Fig 5.A. Based on our experiments in section 4.4, the incorporation of micro pathway in the early epochs of training process overshadows the effect of our macro pathway. Not only does this phenomenon lead to a model that overfits low-level features of the sketch, but it also prevents the model from generalizing to broader spatial and conceptual features. To mitigate this, we employ a modulator that progressively increases the impact of the Micro Pathway, i.e. the FGC module, during training. The modulator is based on a smooth tanh function:
@@
Here, $t$ is the current epoch, $T$ is the total number of epochs of training, $k=6$, $\psi\in[-3,3]$, $m_{\text{min}}=0.2$ and $m_{\text{max}}=1$ where $m_{\text{min}}$ and $m_{\text{max}}$ define the range within which the modulator effect (in percent), i.e. $m_{t}$, will vary over the course of the epochs. In order to choose $m_{\text{min}}$, we heuristically found that the maximum lower bound for negligible effect of the FGC is at $m_{\text{min}}=0.2$. We did not conduct an extensive hyperparameter search for $m_{\text{min}}$ and only chose this value based on our observation of different case studies, Figure 2. As seen in Figure 5.A, the module ensures that diffusion is more affected by the Macro Pathway and less by the Micro Pathway in the early stages of training. As the training progresses, $m_{t}$ for the Micro Pathway approaches 1 and as a result our FGC module will have an equal impact in the training as that of the CGC. By gradually modulating the influence of the Micro Pathway, we prevent the premature weakening of high-level spatial layout presented by the Macro Pathway, and ensure that both pathways contribute optimally throughout the training process. The effectiveness of our modulator is experimented in section 4.4.
@@ 3.4 Inference Knob
In typical diffusion models, the early denoising steps during inference focus on generating high-level spatial features, while the later steps refine finer details (Ho et al., 2020; Meng et al., 2021). In our dual-pathway model, this mechanism is explicitly implemented by our proposed inference-time Knob. This is essentially a user-controlled tool that determines the range of how much abstraction or rigid alignment with respect to the input sketch is desired by the user, Fig 5.C.
@@
[…]
@@
A lower $\gamma$ value results in more abstract outputs with respect to the original input sketch, while a higher value makes the model produce images that closely match the sketch’s finer details. This adaptive control allows KnobGen to accommodate a wide range of user preferences and input complexities, ensuring that both novice and artists can generate images that align with their expectations, Figure 6. The effectiveness of our proposed Knob mechanism is illustrated in Appendix( C.2).
@@ 4 Experiment
The extension of the qualitative experiments is available in the Appendix (C). Furthermore, details about the setup used in the training and evaluation are discussed at length in the Appendix (A).
@@ 4.1 Qualitative Results
Our qualitative results demonstrate the flexibility and effectiveness of KnobGen in handling varying sketch qualities. As shown in Figure 1, KnobGen is able to seamlessly adapt to sketches from rough amateur drawings to refined professional ones, highlighting its ability to cover the entire spectrum of user expertise. Figure 6 illustrates the impact of our knob mechanism, where increasing the knob value (left to right) progressively improves the fidelity to the sketch input. This dynamic adjustment enables precise control over the level of detail, allowing users to fine-tune generation outputs. More qualitative results are provided in the Appendix (C.2).
@@
In order to conduct a fair comparative study, we evaluated KnobGen against baselines such as (Mou et al., 2024; Li et al., 2024; Zhang et al., 2023) on professional-grade sketches, novice ones and a spectrum in between. Figure 4 illustrates the superior quality of the novice-based sketch conditioning using our method against all the other baselines. KnobGen not only captures the spatial layout of the input sketch thanks to the CGC module but also extends beyond it by generating fine-grained details through the FGC module which ultimately produces a naturally appealing images. Whereas the baselines either rigidly conditions themselves on the imperfect input sketch or does not follow the spatial layout desired by the user.
@@
[…]
Table 1: Comparison of various models on CLIP score (higher is better), FID score (lower is better), and Aesthetic score (higher is better). The models include ControlNet (CNet), T2I-Adapter (T2I), UniControl (UC), ControlNet++ (CNet++), AnimateDiff (ADiff), KnobGen with ControlNet as the Fine-Grained Controller (KG-CN), and KnobGen with T2I-Adapter as the Fine-Grained Controller (KG-T2I). KnobGen variants (KG-CN and KG-T2I) consistently outperform other models. The number of sketches used for the evaluation is 600.
@@ 4.3 Quantitative Results
Table 1 provides a quantitative comparison between state-of-the-art DM models and KnobGen. We evaluated our model with two different FGC module plugins, that is, ControlNet and T2I-Adapter. We call our KnobGen whose FGC module is ControlNet KG-CN and with the T2I-Adapter KG-T2I. We measure performance using the CLIP score (prompt-image alignment), Fréchet Inception Distance (FID) and Aesthetic score (for more information, see Appendix A). KG-CN achieves the highest CLIP score of 0.3353, surpassing the best baseline of 0.3214. KG-CN also gives the lowest FID score (93.87) and the highest aesthetic score (0.5349), demonstrating superior image quality and realism. We use a stratified sampling method based on pixel count to evaluate professional and amateur sketches, ensuring robustness across varying complexity levels. Our results demonstrate KnobGen’s effectiveness in generating high-quality images, regardless of input skill level.
@@ 4.4 Ablation Study
Figure 7. presents the results of these experiments, showcasing images generated with and without the Modulator. The comparative analysis reveals that the model trained with the Modulator exhibits a significantly enhanced ability to integrate sketch-based coarse-grained guidance into the image generation process. This indicates that the Modulator not only improves the model’s overall performance but also ensures that the CGC’s influence is effectively optimized during training, resulting in higher-quality, more accurate image synthesis.
@@ 5 Conclusion
Reproducibility Statement
Our experiment setups are concisely described in Section 4, with additional implementation details provided in Appendix Sections A and B. We will make our code publicly available to facilitate the reproduction of our results upon acceptance of this paper.
@@
  • Appendix A: details the setup of training and evaluation in KnobGen.
  • Appendix B: expands on the details of the proposed CGC module.
  • Appendix C: provides more qualitative results.
Appendix A Setup
@@
We utilized the MultiGen-20M dataset, as introduced by Qin et al. (2023), to train and evaluate our model. The dataset offers various conditions, making it a suitable choice for our approach. We selected 20,000 images for training, focusing specifically on those with the Holistically-nested Edge Detection (HED) (Xie & Tu, 2015) condition. However, we modified the KnobGen condition by applying a thresholding technique, where pixels below a threshold value of 50 were set to zero, and those above were set to one. This threshold value was chosen through simple visual comparisons of several samples using different thresholds, allowing us to identify the most effective value. This modification essentially transforms the HED condition into a sketch. For evaluation, we curated two distinct sets of images. The first evaluation set consisted of 500 randomly selected samples, which are similar to a sketch drawn by a seasoned artist (we followed the thresholding technique for this part), allowing us to measure our model’s effectiveness in professional settings. To further test the robustness and adaptability of our approach, we compiled a second evaluation set of 100 hand-drawn images created by non-professional individuals. This diverse testing set enabled us to demonstrate the model’s ability to generalize across a broad spectrum of users, ensuring it can handle both professionally designed and amateur drawings with high robustness.
@@
In this work, we evaluate the performance of our proposed model against several state-of-the-art (SOTA) diffusion-based models. Specifically, we conduct both qualitative and quantitative comparisons with prominent models such as ControlNet (Zhang et al., 2023), T2I-Adapter (Mou et al., 2024), AnimateDiff (Guo et al., 2023), UniControl (Qin et al., 2023), and ControlNet++ (Li et al., 2024). These models have achieved significant advances in fine-grained control of image generation by incorporating sketch-based conditions into the diffusion process. Since AnimateDiff is a video-based DM, we only use the first frame of the generated video by it as the comparison point.
@@
We perform qualitative and quantitative evaluation. In the qualitative evaluation, we compare our model’s performance across different scenarios of varying input conditions and complexities. For quantitative evaluation, we utilize several metrics to assess the quality of the generated images. First, we calculate the Fréchet Inception Distance (FID) (Heusel et al., 2017; Karras et al., 2019), which measures the similarity between generated and natural images using a pre-trained InceptionV3 model (Szegedy et al., 2016). Lower FID values indicate better generation quality; we used Hukkelås (2020) implementation for our evaluation, which used the default pre-trained InceptionV3 model available in Pytorch (Paszke et al., 2017). To evaluate the alignment between the generated images and the text prompts, we use CLIP (Radford et al., 2021), specifically the pre-trained DetailCLIP model (Monsefi et al., 2024) with a Vision Transformer (ViT-B/16) backbone. Higher CLIP scores signify better alignment between the generated images and their corresponding prompts. Finally, we assess the realism and aesthetic quality of the generated images using the metric proposed by (Ke et al., 2023), where higher scores reflect more realistic and visually appealing images.
@@
Our proposed KnobGen framework is built on top of Stable Diffusion v1.5 (Rombach et al., 2022), with the original parameters kept frozen throughout training. For the Fine-Grained Controller (FGC) module, we employed two different pre-trained models to demonstrate the flexibility and effectiveness of our approach across multiple setups. Specifically, we integrated ControlNet (Zhang et al., 2023) and T2I-Adapter (Mou et al., 2024), both of which had their parameters frozen and were not updated during training. The architecture and integration of these components are illustrated in Figure 5. We trained the CGC module for a total of $2000$ epochs using 16 A100 GPUs. During the initial $1500$ epochs, we employed the modulator mechanism, as described in Section 3.3, with a learning rate of $51e-$ 5. In the final $500$ epochs, we fine-tuned the CGC model with a reduced learning rate of $61e-$ 6 to ensure robustness and to improve the quality of the generated images.
Appendix B Model architecture
@@
Appendix C More Qualitative Result

### v3

@@ 3.2.1 Macro Pathway
Diffusion models typically rely on text-based conditioning using CLIP text encoders [32] to capture high-level semantics [34, 38, 26], but this approach often misses out on structural cues inherent to other modalities, such as sketches. Although models such as CLIP [32] encode visual features and textual semantics, they remain biased toward coarse-grained features [1, 48]. In our CGC module, Figure 5.B, we used this fact to our advantage to fuse a high-level visual and linguistic understanding to control DM generation by incorporating both text and image embeddings through a cross-attention mechanisms.
@@
In our CGC module, we leverage the trained CLIP text encoder and its corresponding image encoder variant available in the pretrained Stable Diffusion Model [36]. Our CGC module first takes the raw sketch image (condition) and prompt as input. Using the CLIP image and text encoders, the CGC module first projects them into $x_{i}\in\mathbb{R}^{256\times 1024}$ and $x_{p}\in\mathbb{R}^{77\times 768}$ which are the image and text embeddings. A cross-attention mechanism then fuses these embeddings to produce a multimodal representation that combines textual semantics and visual cues. This enables the diffusion process to encode the high-level semantics from text while explicitly integrating spatial features from the sketch using the Clip image encoder. The cross-attended embeddings are injected into layers of the denoising U-Net to preserve the coarse-grained visual-textual features throughout the diffusion process. Detailed discussion of the CGC module is in the Appendix 7.
@@ 3.2.2 Micro Pathway
For artistic users, preserving fine-grained details such as object boundaries and textures is essential. The Fine-Grained Controller (FGC) is designed to address these requirements by integrating pretrained modules such as ControlNet [54] and the T2I-Adapter [24], which excel in capturing these intricate features. Our Micro Pathway can utilize any pretrained fine-grained controller module which shows the flexibility of our proposed framework.
@@ 3.3 Modulator at training
One of the key innovations in KnobGen is the tanh-based modulator, which regulates the contributions of the micro and macro pathways during training, Figure 5.A. Based on our experiments in section 4.4, the incorporation of micro pathway in the early epochs of training process overshadows the effect of our macro pathway. Not only does this phenomenon lead to a model that overfits low-level features of the sketch, but it also prevents the model from generalizing to broader spatial and conceptual features. To mitigate this, we employ a modulator that progressively increases the impact of the Micro Pathway, i.e. the FGC module, during training. The modulator is based on a smooth tanh function:
@@
Here, $t$ is the current epoch, $T$ is the total number of epochs of training, $k=6$, $\psi\in[-3,3]$, $m_{\text{min}}=0.2$ and $m_{\text{max}}=1$ where $m_{\text{min}}$ and $m_{\text{max}}$ define the range within which the modulator effect (in percent), i.e. $m_{t}$, will vary over the course of the epochs. In order to choose $m_{\text{min}}$, we heuristically found that the maximum lower bound for negligible effect of the FGC is at $m_{\text{min}}=0.2$. We did not conduct an extensive hyperparameter search for $m_{\text{min}}$ and only chose this value based on our observation of different case studies. As seen in Figure 5.A, the module ensures that diffusion is more affected by the Macro Pathway and less by the Micro Pathway in the early stages of training. As the training progresses, $m_{t}$ for the Micro Pathway approaches 1 and as a result our FGC module will have an equal impact in the training as that of the CGC. By gradually modulating the influence of the Micro Pathway, we prevent the premature weakening of high-level features presented by the Macro Pathway, and ensure that both pathways contribute optimally throughout the training. The effectiveness of our modulator is experimented in section 4.4.
Remark.We selected the tanh function for its gradual transition across epochs which enables balanced modulation of coarse and fine-grained contributions during training. While we did not test other functions, the tanh function’s properties effectively support stable learning.
@@ 3.4 Inference Knob
In typical diffusion models, the early denoising steps during inference focus on generating high-level spatial features, while the later steps refine finer details [11, 21]. In our dual-pathway model, this mechanism is explicitly implemented by our proposed inference-time Knob. This is essentially a user-controlled tool (Figure 5.C) that determines the range of how much abstraction or rigid alignment with respect to the input sketch is desired by the user.
@@
$f_{\ell}(t)=\begin{cases}f_{\text{coarse}}(t)+f_{\text{fine}}(t),&\text{if }t\leq\gamma,\\ f_{\text{coarse}}(t),&\text{if }t>\gamma,\end{cases}\quad\begin{subarray}{c}\forall\,\ell\in\\ \{\text{U-Net layers}\}\end{subarray}$
@@
A lower $\gamma$ value results in more abstract outputs with respect to the original input sketch, while a higher value makes the model produce images that closely match the sketch’s finer details. This adaptive control allows KnobGen to accommodate a wide range of user preferences and input complexities shown in Figure 7, ensuring that both novice and artists can generate images that align with their expectations. The effectiveness of our proposed Knob mechanism is illustrated in Appendix( 9.2).
Remark. Unlike the T2I-Adapter [24] weighting approach, which applies a uniform weight over the entire denoising process (see Figure 2), our knob mechanism introduces a flexible adjustment. This mechanism allows users to selectively balance fine-grained and coarse-grained details throughout denoising, tailoring image generation to the preferred level of detail, as demonstrated in Figure 7.
@@ 4 Experiment
The extension of the qualitative experiments is available in the Appendix (9). More quantitative and user study result are in Appendix (8). Furthermore, details about the setup used in the training and evaluation are in the Appendix (6).
@@ 4.1 Qualitative Results
Our qualitative results demonstrate the flexibility and effectiveness of KnobGen in handling varying sketch qualities. KnobGen is able to seamlessly adapt to sketches from rough amateur drawings to refined professional ones, highlighting its ability to cover the entire spectrum of user expertise. Figure 7 illustrates the impact of our knob mechanism, where increasing the knob value (left to right) progressively improves the fidelity to the sketch input. This dynamic adjustment enables precise control over the level of detail, allowing users to fine-tune generation outputs. More qualitative results with different input conditions and modes, such as no prompt, professional sketch and free-chyle sketch are provided in the Appendix (9.2).
@@
In order to conduct a fair comparative study, we evaluated KnobGen against baselines such as [24, 19, 54] on professional-grade sketches, novice ones and a spectrum in between. Figure 4 illustrates the superior quality of the novice-based sketch conditioning using our method against all the other baselines. KnobGen not only captures the spatial layout of the input sketch thanks to the CGC module but also extends beyond it by generating fine-grained details through the FGC module which ultimately produces a naturally appealing images. Whereas the baselines either rigidly conditions themselves on the imperfect input sketch or does not follow the spatial layout desired by the user.
@@
CLIP $\uparrow$ 0.3214 0.3152 0.3210 0.3204 0.2988 0.3353 0.3271
FID $\downarrow$ 106.25 109.75 95.30 99.51 119.01 93.87 98.41
Aesthetic $\uparrow$ 0.5182 0.5093 0.5133 0.5253 0.4751 0.5349 0.5208
Table 1: Model comparison on CLIP, FID, and Aesthetic scores. Models include ControlNet (CNet), T2I-Adapter (T2I), UniControl (UC), ControlNet++ (CNet++), AnimateDiff (ADiff), and KnobGen variants (KG-CN, KG-T2I) with ControlNet and T2I-Adapter as Fine-Grained Controllers, respectively. KnobGen variants consistently outperform other models.
@@ 4.3 Quantitative Results
Table 1 provides a quantitative comparison between state-of-the-art DM models and KnobGen over 600 sketch images. We evaluated our model with two different FGC module plugins, that is, ControlNet and T2I-Adapter. We call our KnobGen whose FGC module is ControlNet KG-CN and with the T2I-Adapter KG-T2I. We measure performance using the CLIP score (prompt-image alignment), Fréchet Inception Distance (FID) and Aesthetic score (for more information, see Appendix 6). KG-CN achieves the highest CLIP score of 0.3353, surpassing the best baseline of 0.3214. KG-CN also gives the lowest FID score (93.87) and the highest aesthetic score (0.5349), demonstrating superior image quality and realism. We use a stratified sampling method based on pixel count to evaluate professional and amateur sketches, ensuring robustness across varying complexity levels. Our results demonstrate KnobGen’s effectiveness in generating high-quality images, regardless of input skill level.
@@ 4.4 Ablation Study
Figure 6. presents the results of these experiments, showcasing images generated with and without the Modulator. The comparative analysis reveals that the model trained with the Modulator exhibits a significantly enhanced ability to integrate sketch-based coarse-grained guidance into the image generation process. This indicates that the Modulator not only improves the model’s overall performance but also ensures that the CGC’s influence is effectively optimized during training, resulting in controlled image synthesis.
@@ 5 Conclusion
Supplementary Material
@@
  • Appendix 6: explain training and evaluation setup.
  • Appendix 7: expands on the details of the CGC module.
  • Appendix 8: explains our conducted user study.
  • Appendix 9: provides more qualitative results.
6 Setup
@@
We utilized the MultiGen-20M dataset, as introduced by [31], to train and evaluate our model. The dataset offers various conditions, making it a suitable choice for our approach. We selected 20,000 images for training, focusing specifically on those with the Holistically-nested Edge Detection (HED) [49] condition. However, we modified the KnobGen condition by applying a thresholding technique, where pixels below a threshold value of 50 were set to zero, and those above were set to one. This threshold value was chosen through simple visual comparisons of several samples using different thresholds, allowing us to identify the most effective value. This modification essentially transforms the HED condition into a sketch. For evaluation, we curated two distinct sets of images. The first evaluation set consisted of 500 randomly selected samples, which are similar to a sketch drawn by a seasoned artist (we followed the thresholding technique for this part), allowing us to measure our model’s effectiveness in professional settings. To further test the robustness and adaptability of our approach, we compiled a second evaluation set of 100 hand-drawn images created by non-professional individuals. This diverse testing set enabled us to demonstrate the model’s ability to generalize across a broad spectrum of users, ensuring it can handle both professionally designed and amateur drawings with high robustness.
@@
In this work, we evaluate the performance of our proposed model against several state-of-the-art (SOTA) diffusion-based models. Specifically, we conduct both qualitative and quantitative comparisons with prominent models such as ControlNet [54], T2I-Adapter [24], AnimateDiff [8], UniControl [31], and ControlNet++ [19]. These models have achieved significant advances in fine-grained control of image generation by incorporating sketch-based conditions into the diffusion process. Since AnimateDiff is a video-based DM, we only use the first frame of the generated video by it as the comparison point.
@@
We perform qualitative and quantitative evaluation. In the qualitative evaluation, we compare our model’s performance across different scenarios of varying input conditions and complexities. For quantitative evaluation, we utilize several metrics to assess the quality of the generated images. First, we calculate the Fréchet Inception Distance (FID) [9, 15], which measures the similarity between generated and natural images using a pre-trained InceptionV3 model [45]. Lower FID values indicate better generation quality; we used [12] implementation for our evaluation, which used the default pre-trained InceptionV3 model available in Pytorch [27]. To evaluate the alignment between the generated images and the text prompts, we use CLIP [32], specifically the pre-trained DetailCLIP model [22] with a Vision Transformer (ViT-B/16) backbone. Higher CLIP scores signify better alignment between the generated images and their corresponding prompts. Finally, we assess the realism and aesthetic quality of the generated images using the metric proposed by [17], where higher scores reflect more visually appealing images.
@@
Our proposed KnobGen framework is built on top of Stable Diffusion v1.5 [36], with the original parameters kept frozen throughout training. For the Fine-Grained Controller (FGC) module, we employed two different pre-trained models to demonstrate the flexibility and effectiveness of our approach across multiple setups. Specifically, we integrated ControlNet [54] and T2I-Adapter [24], both of which had their parameters frozen and were not updated during training. The architecture and integration of these components are illustrated in Figure 5. We trained the CGC module for a total of $2000$ epochs using 16 A100 GPUs. During the initial $1500$ epochs, we employed the modulator mechanism, as described in Section 3.3, with a learning rate of $51e-$ 5. In the final $500$ epochs, we fine-tuned the CGC model with a reduced learning rate of $61e-$ 6 to ensure robustness and to improve the quality of the generated images.
7 Model architecture
@@
8 More Quantitative Result: User Study
In this section we provide an additional quantitative evaluation which is our conducted user-study experiment.
Figure 9: Comparison of model performance in our user study: Our user study includes three metrics sketch alignment, prompt alignment, and aesthetic score. While all models perform similarly in prompt alignment, KnobGen significantly outperforms ControlNet and T2I-Adapter in sketch alignment and aesthetic score, demonstrating its superior capability in handling sketch precision and generating visually appealing outputs.
We conducted a user study to gain deeper insights into the perceived quality and usability of KnobGen, compared to existing state-of-the-art baselines. This study involved 100 participants, who were asked to evaluate a set of 10 images selected randomly from a set of 50 generated images with different complexities across three key dimensions: sketch alignment, prompt alignment, and aesthetic quality. Sketch alignment measures how well the generated image adheres to the spatial and structural details of the input sketch, while prompt alignment assesses the consistency between the generated image and the provided textual prompt. Aesthetic score captures the overall visual appeal of the generated images. Participants rated each dimension on a scale from 0 to 10, with higher scores indicating better performance. While models are equally performing in the prompt alignment as shown in Figure 9, KnobGen exhibits a distinct edge over handling diverse set of sketches while maintaining a visually appealing generation.
9 More Qualitative Result